OpenStudy (anonymous):

A projectile is fired with an initial speed of 38.8 m/s. Find the angle of projection such that the maximum height of the projectile is equal to its horizontal range.

OpenStudy (anonymous):

sorry...initial speed is 38.8

OpenStudy (jt950):

Uy=USinX Ux=UCosX (X is the angle of projection) Flight time, T=2t (T=total flight time,t is time taken to maximum=time taken from maximum to ground) V=Uy+gt 0=USinx+gt 2t=2USinx/g T=2USinx/g Maximum Height , V=0,U=38.8m/s V=0 V^2=Uy^2+2gS 0=Uy^2+2gS S=(U^2)(sinx^2)/2g Range, R=(Vx)(T) =(UCosx)(2UsinX/g) =(U^2)(Sin2x)/g Maximum height = (u^2)(sinx^2)/2g Range = (u^2)(sin2x)/g Take Maximum height = Range ( u=38.8m/s, g = 9.8m/s) you can find x the angle of projection Work it out yourself , I hope this help.

OpenStudy (jt950):